# Architecture-Aware Learning Curve Extrapolation via Graph Ordinary Differential Equation

### v1

@@ I Introduction
Training neural architectures is a resource-intensive endeavor, often demanding considerable computational power and time. Researchers have developed various methodologies to predict the performance of neural networks early in the training process using learning curve data. Some methods Domhan et al. (2015); Gargiani et al. (2019); Adriaensen et al. (2023) apply Bayesian inference to project these curves forward, while others employ time-series prediction techniques, such as LSTM networks. Despite their effectiveness, these approaches (Swersky et al., 2014; Baker et al., 2017) typically overlook the architectural features of networks, missing out on crucial insights that could be derived from the models’ topology.
@@
On another front, architecture-based predictive models have been developed to forecast network performance based purely on NN structures Shi et al. (2019); Friede et al. (2019); Wen et al. (2020); Tang et al. (2020); Yan et al. (2020); Ning et al. (2020); Xu et al. (2019); Siems et al. (2020). These models facilitate a deeper understanding of the relationship between architectures and their performance. However, they are limited in their ability to predict precise learning curve values at specific epochs and struggle to capture the variability in performance that a single architecture can exhibit under diverse training conditions.
@@
Moreover, there is a growing interest in conceptualizing the optimization process during NN training as a dynamical system. By considering the step size in gradient descent as approaching zero, it is possible to formulate an ordinary differential equation for the model parameters Su et al. (2016); Zhang et al. (2024); Maskan et al. (2023). This perspective is useful for analyzing the convergence of different optimization algorithms, especially for convex problems. Building on this foundational idea, we propose an innovative approach that models the evolution of learning curves using neural differential equations Chen et al. (2018), tailored for a inductive setting where the trained learning curve predictor is applicable to new learning curves generated under various training configurations, such as different architectures, batch sizes, and learning rates. This approach leverages recent advancements in differential equations to provide a flexible framework capable of handling the complexities of modern neural training processes.
@@
Our method merges the structural attributes of neural architectures with the dynamic nature of learning curves. We utilize a seq2seq variational autoencoder framework to analyze the initial stages of a learning curve and predict its future progression. This predictive capability is further enhanced by an architecture-aware component that produces a graph-level embedding from the architecture’s topology, employing techniques like Graph Convolutional Networks (GCN) Kipf and Welling (2016) and Differentiable Pooling Ying et al. (2018). This integration not only improves the accuracy of learning curve extrapolations compared to existing methods but also significantly facilitates model ranking, potentially leading to more efficient use of computational resources, accelerated experimentation cycles, and faster progress in the field of machine learning.
@@ II Related Work
Previous studies have explored learning curve prediction through diverse approaches Swersky et al. (2014); Domhan et al. (2015); Baker et al. (2017); Chandrashekaran and Lane (2017); Gargiani et al. (2019); Ru et al. (2021); Klein et al. (2022); Adriaensen et al. (2023). A line of work has focused on Bayesian frameworks. Specifically, Domhan et al. (2015) utilized a weighted combination of functions to predict mean future validation accuracy and facilitate early termination of underperforming training runs. Building on this, Chandrashekaran and Lane (2017) extended basis function extrapolation by incorporating historical learning curves from previous training runs, while Klein et al. (2022) proposed a Bayesian neural network to flexibly model learning curves, removing the constraint that each epoch must outperform the previous one and thereby reducing instability in predictions. More recently, Adriaensen et al. (2023) applied a prior-data-fitted network training paradigm to enhance sampling efficiency from the posterior distribution of learning curves. Despite these advancements, existing methods overlook the role of architectural design, whereas our work explicitly incorporates this information to better model and understand the evolution of learning curves.
Architecture-based performance prediction. Advances in Graph Neural Networks (GNNs) have led to innovative methods for predicting the performance of neural network architectures Shi et al. (2019); Friede et al. (2019); Wen et al. (2020); Tang et al. (2020); Yan et al. (2020); Ning et al. (2020); Xu et al. (2019); Siems et al. (2020). In exploring unsupervised learning strategies, Yan et al. Yan et al. (2020) used architecture embeddings created by a pre-trained model to feed a Gaussian Process model for performance prediction. These studies often incorporate foundational GNN models such as the GCN Kipf and Welling (2016) and the Graph Isomorphism Network (GIN) Xu et al. (2018) to effectively process input architectures. Various training objectives have been considered, including Mean Squared Error (MSE) Shi et al. (2019); Wen et al. (2020); Ning et al. (2020); Tang et al. (2020), graph reconstruction loss Friede et al. (2019); Tang et al. (2020); Yan et al. (2020), and pair-wise ranking loss Ning et al. (2020); Xu et al. (2019), highlights the diverse methods aimed at improving the prediction of architecture performance. Building on these foundations, our approach goes beyond simply using the graph representation to extract a scalar performance value; instead, it integrates the graph information into the ODE of loss and can extrapolate to any time step of interest, including the value at convergence.
@@
Neural ordinary differential equations (NODE) Chen et al. (2018) introduce a general framework for parameterizing ODEs with deep neural networks, deriving backpropagation through the adjoint sensitivity method. Variational autoencoders combined with NODE, as introduced in Rubanova et al. (2019), are used to predict dynamics from irregularly sampled time-series data. Building on these developments, recent works Huang et al. (2020, 2021); Luo et al. (2023); Huang et al. (2024a); Jiang et al. (2023); Huang et al. (2024b) further integrate graph neural networks (GNNs) with NODE to model temporal graphs, allowing for the representation of evolving nodes and edges over time. In contrast to modeling individual nodal trajectories, this work employs GNNs as a graph reduction technique, using their embeddings to drive the evolution of learning curves within the NODE framework.
@@ III-A Problem Formulation
Learning curves, such as train or test loss, are generated by optimizing a neural network on various source tasks including adult income classification, image classification, and housing price regression Vanschoren et al. (2014). Our goal is to train a single latent ODE model capable of extrapolating learning curves for a given source task across different architectures. The model infers full learning curves of length $m$ using only the initial $n$ epochs $y_{1},\ldots,y_{n}$ and the corresponding network architecture, denoted as $G$.
@@
We employ a seq2seq variational autoencoder framework Rubanova et al. (2019); Huang et al. (2020, 2021), where a sequence encoder parameterized by $\phi$ processes the early part of the learning curve to estimate the variational parameters of the posterior distribution $q_{\phi}$, which determines the latent state at the start of the prediction period, represented by $\boldsymbol{z}_{n+1}\in\mathbb{R}^{D}$. A numerical solver then integrates an ODE function, denoted as $f:\mathbb{R}^{D}\rightarrow\mathbb{R}^{D}$, governing this latent state, starting from the initial condition $\boldsymbol{z}_{n+1}$. Each subsequent latent state $\boldsymbol{z}_{i}$ is decoded independently to predict the output $\hat{y}_{i}$ for $i>n$. This process is mathematically represented as follows:
@@ III-B Architecture-aware Differential Equation
We implement this using an RNN with GRU units for the sequence encoder. Other encoder options include Self-Attention Vaswani et al. (2017) and Temporal Convolutional Networks (TCN) Pandey and Wang (2019), which adapt well to varying observation lengths. We conduct an ablation study in the experimental section to evaluate the performance of these alternative sequence encoder implementations.
@@
Assuming the graph representation $G$ contains $N$ nodes, the adjacency matrix $A\in\mathbb{R}_{\geq 0}^{N\times N}$ details node connections, where $A_{ij}>0$ represents the edge from node $i$ to node $j$. We analyze two foundational neural network types: MLPs and CNNs. In MLPs, nodes correspond to neurons and edges correspond to the presence of the connections between neurons, whereas in CNNs, nodes represent feature maps and edges depict operations like $1\times 1$ and $3\times 3$ convolutions, or $3\times 3$ average pooling. For CNNs, we adopt the cell-based representation Dong and Yang (2020); Liu et al. (2018), which consists of four principal building blocks: stems, normal cells, reduction cells, and classification heads. The stem block is a fixed sequence of convolutional layers to process the input images. This is followed typically by 14-20 cells with reduction cells placed at 1/3 and 2/3 of the total depth. A normal cell contains 4 nodes, each of which belongs to the set of operations: skip connections, identity or zero (indicating the presence or absence of connections between certain layers), $1\times 1$ and $3\times 3$ convolutions, and $3\times 3$ average pooling. Finally, the classification head employs a global pooling layer followed by a single fully connected layer and returns the network’s output. Since the cell is repeated throughout this macro-skeleton, a CNN can be uniquely represented by its cell.
To derive a graph-level embedding, we first implement node-level message passing, followed by global pooling to extract a global embedding. Our method differs from existing architecture-based performance prediction approaches Liu et al. (2018); Wen et al. (2020); Knyazev et al. (2021) by treating the node as a feature map rather than focusing on operations. Nodal features are calculated from the in-degree and out-degree of each node, normalized by the total number of edges. For MLPs, edge features are binary, while for CNN cells, they are integers representing operation types: $\{\text{zeroize}:0,\text{$1\times 1$ conv}:1,\text{$3\times 3$ conv}:2,\text{$3\times 3$ avg pooling}:3\}$. The nodal feature matrix is denoted as $X\in\mathbb{R}^{N\times d}$. The graph encoding process returns a vector representation of the architecture.
@@
For node-level message passing, we employ GCN layers and normalize the adjacency matrix to stabilize training, following  Kipf and Welling (2016).
@@ A GCN layer then transforms the nodal features into a hidden representation:
The transformation from discrete to continuous domain is predicated on the assumption that the step size, or learning rate, approaches zero, thereby approximating the time derivative of NN parameters $\Theta$ using $\frac{d\Theta}{dt}=-\frac{\partial\mathcal{L}}{\partial\Theta}$ Su et al. (2016), where $\mathcal{L}$, $\Theta$ denotes the objective function and parameters from the source task. Assuming the parameters implicitly depends on time $t$, the time derivative of the training loss can be written as:
@@
Here, $\dot{\boldsymbol{z}}$ denotes the derivative of the latent state vector $\boldsymbol{z}$ with respect to time, driven by the function $f_{\theta_{2}}$, which takes as input both the latent state $\boldsymbol{z}$ and a graph-level embedding $\boldsymbol{z}_{G}$. This combination allows the model to simultaneously consider the dynamic properties of the learning curve and the static characteristics of the architecture, enhancing the predictive capability of the system. Eq (9) can be regarded as a single-agent representation of the dynamics induced by NN training, which involves multiple trajectories of neurons, edge weights, and loss. This reduced representation of coupled dynamical system has been explored in Gao et al. (2016); Laurence et al. (2019) to study the tipping point of the original network dynamics. The difference from the prior dynamical system reduction approach is that the graph reduction mechanism is learnable so that the model can be adapted to unseen trajectories derived from different optimization trials.
@@ IV Experiments
We consider test loss and test accuracy curves for both MLP-based and CNN-based architectures. Specifically for MLPs, we use car and segment tabular data binary classification tasks from OpenML Vanschoren et al. (2014) as source tasks. Following LCBench Zimmer et al. (2021), we randomly generate training configurations that include variables such as the number of layers, number of hidden units, and learning rates. For each dataset, we conduct 550 optimization trials across 200 epochs. For CNN-based models, we employ the NAS-Bench-201 dataset Dong and Yang (2020), which provides comprehensive learning curves for each architecture over a span of 200 epochs across two image datasets: CIFAR-10, CIFAR-100 Krizhevsky et al. (2009). We randomly select 5,000 architectures from CIFAR-10 and CIFAR-100 to form our dataset. Furthermore, we reserve 20% of all trials as the test set for each MLP source task and 25% for each CNN source task.
@@ IV-A Extrapolating Real-world Learning Curves
The goal of this experiment is to evaluate the LC-GODE model against established learning curve prediction methods using real-world benchmarks. We train our model separately on the test loss curves of each source task. The condition length is set to 10 epochs for all methods in this experiment. The instantiation of LC-GODE that we report features: (i) an architecture encoder that utilizes 2 layers and employs a learnable pooling technique, (ii) an observed time-series encoder implemented using GRU, (iii) an ODE function with a 2-layer MLP and integrated using the Runge-Kutta 4 method Butcher (1996). Further details on the evaluation metrics and the training settings can be found in the supplemental materials.
@@
We evaluate our model against six methods, including three Bayesian approaches and three general time-series prediction approaches. (i) LC-BNN: This method utilizes a Bayesian neural network to model the posterior distribution of future learning curves. The probability function is constructed from a combination of basis functions, following the approach described by Domhan et al. Domhan et al. (2015). (ii) LC-PFN: A transformer-based model is trained on synthetic curves that are generated from a pre-defined prior. This method serves as an efficient alternative to traditional Markov Chain Monte-Carlo (MCMC) techniques for sampling from posterior distributions. (iii) VRNN: A probabilistic model utilizing random forests and Bayesian recurrent neural networks. (iv) LSTM: This Recurrent Neural Network captures sequential dependencies. Before the observation cutoff at $n$, input states are derived from actual observations. Post $n$, the model uses its own predictions from previous timestamps as inputs. (v) NODE: Latent Ordinary Differential Equations, focusing primarily on modeling the latent loss representation without incorporating architectural information. (vi) NSDE: Latent Stochastic Differential Equations, consisting of a drift term and a diffusion term. The drift term is the same as NODE, and the sequence encoder and decoder is the same as in our model. Both NODE and NSDE are trained using the variational framework.

### v3

@@ I Introduction
Training neural architectures is a resource-intensive endeavor, often demanding considerable computational power and time. Researchers have developed various methodologies to predict the performance of neural networks early in the training process using learning curve data. Some methods (Domhan et al., 2015; Gargiani et al., 2019; Adriaensen et al., 2023) apply Bayesian inference to project these curves forward, while others employ time-series prediction techniques, such as LSTM networks. Despite their effectiveness, these approaches (Swersky et al., 2014; Baker et al., 2017) typically overlook the architectural features of networks, missing out on crucial insights that could be derived from the models’ topology.
@@
On another front, architecture-based predictive models have been developed to forecast network performance based purely on NN structures (Shi et al., 2019; Friede et al., 2019; Wen et al., 2020; Tang et al., 2020; Yan et al., 2020; Ning et al., 2020; Xu et al., 2019; Siems et al., 2020). These models facilitate a deeper understanding of the relationship between architectures and their performance. However, they are limited in their ability to predict precise learning curve values at specific epochs and struggle to capture the variability in performance that a single architecture can exhibit under diverse training conditions.
@@
Moreover, there is a growing interest in conceptualizing the optimization process during NN training as a dynamical system. By considering the step size in gradient descent as approaching zero, it is possible to formulate an ordinary differential equation for the model parameters (Su et al., 2016; Zhang et al., 2024; Maskan et al., 2023). This perspective is useful for analyzing the convergence of different optimization algorithms, especially for convex problems. Building on this foundational idea, we propose an innovative approach that models the evolution of learning curves using neural differential equations (Chen et al., 2018), tailored for a inductive setting where the trained learning curve predictor is applicable to new learning curves generated under various training configurations, such as different architectures, batch sizes, and learning rates. This approach leverages recent advancements in differential equations to provide a flexible framework capable of handling the complexities of modern neural training processes.
@@
Our method merges the structural attributes of neural architectures with the dynamic nature of learning curves. We utilize a seq2seq variational autoencoder framework to analyze the initial stages of a learning curve and predict its future progression. This predictive capability is further enhanced by an architecture-aware component that produces a graph-level embedding from the architecture’s topology, employing techniques like Graph Convolutional Networks (GCN) (Kipf and Welling, 2016) and Differentiable Pooling (Ying et al., 2018). This integration not only improves the accuracy of learning curve extrapolations compared to existing methods but also significantly facilitates model ranking, potentially leading to more efficient use of computational resources, accelerated experimentation cycles, and faster progress in the field of machine learning.
@@ II Related Work
Previous studies have explored learning curve prediction through diverse approaches (Swersky et al., 2014; Domhan et al., 2015; Baker et al., 2017; Chandrashekaran and Lane, 2017; Gargiani et al., 2019; Ru et al., 2021; Klein et al., 2022; Adriaensen et al., 2023). A line of work has focused on Bayesian frameworks. Specifically, Domhan et al. (2015) utilized a weighted combination of functions to predict mean future validation accuracy and facilitate early termination of underperforming training runs. Building on this, Chandrashekaran and Lane (2017) extended basis function extrapolation by incorporating historical learning curves from previous training runs, while Klein et al. (2022) proposed a Bayesian neural network to flexibly model learning curves, removing the constraint that each epoch must outperform the previous one and thereby reducing instability in predictions. More recently, Adriaensen et al. (2023) applied a prior-data-fitted network training paradigm to enhance sampling efficiency from the posterior distribution of learning curves. Despite these advancements, existing methods overlook the role of architectural design, whereas our work explicitly incorporates this information to better model and understand the evolution of learning curves.
Architecture-based performance prediction. Advances in Graph Neural Networks (GNNs) have led to innovative methods for predicting the performance of neural network architectures (Shi et al., 2019; Friede et al., 2019; Wen et al., 2020; Tang et al., 2020; Yan et al., 2020; Ning et al., 2020; Xu et al., 2019; Siems et al., 2020). In exploring unsupervised learning strategies, Yan et al. (2020) used architecture embeddings created by a pre-trained model to feed a Gaussian Process model for performance prediction. These studies often incorporate foundational GNN models such as the GCN (Kipf and Welling, 2016) and the Graph Isomorphism Network (GIN) (Xu et al., 2018) to effectively process input architectures. Various training objectives have been considered, including Mean Squared Error (MSE) (Shi et al., 2019; Wen et al., 2020; Ning et al., 2020; Tang et al., 2020), graph reconstruction loss (Friede et al., 2019; Tang et al., 2020; Yan et al., 2020), and pair-wise ranking loss (Ning et al., 2020; Xu et al., 2019), highlights the diverse methods aimed at improving the prediction of architecture performance. Building on these foundations, our approach goes beyond simply using the graph representation to extract a scalar performance value; instead, it integrates the graph information into the ODE of loss and can extrapolate to any time step of interest, including the value at convergence.
@@
Neural ordinary differential equations (NODE) (Chen et al., 2018) introduce a general framework for parameterizing ODEs with deep neural networks, deriving backpropagation through the adjoint sensitivity method. Variational autoencoders combined with NODE, as introduced in (Rubanova et al., 2019), are used to predict dynamics from irregularly sampled time-series data. Building on these developments, recent works (Huang et al., 2020, 2021; Luo et al., 2023; Huang et al., 2024a; Jiang et al., 2023; Huang et al., 2024b) further integrate graph neural networks (GNNs) with NODE to model temporal graphs, allowing for the representation of evolving nodes and edges over time. In contrast to modeling individual nodal trajectories, this work employs GNNs as a graph reduction technique, using their embeddings to drive the evolution of learning curves within the NODE framework.
@@ III-A Problem Formulation
Learning curves, such as train or test loss, are generated by optimizing a neural network on various source tasks including adult income classification, image classification, and housing price regression (Vanschoren et al., 2014). Our goal is to train a single latent ODE model capable of extrapolating learning curves for a given source task across different architectures. The model infers full learning curves of length $m$ using only the initial $n$ epochs $y_{1},\ldots,y_{n}$ and the corresponding network architecture, denoted as $G$.
@@
We employ a seq2seq variational autoencoder framework (Rubanova et al., 2019; Huang et al., 2020, 2021), where a sequence encoder parameterized by $\phi$ processes the early part of the learning curve to estimate the variational parameters of the posterior distribution $q_{\phi}$, which determines the latent state at the start of the prediction period, represented by $\boldsymbol{z}_{n+1}\in\mathbb{R}^{D}$. A numerical solver then integrates an ODE function, denoted as $f:\mathbb{R}^{D}\rightarrow\mathbb{R}^{D}$, governing this latent state, starting from the initial condition $\boldsymbol{z}_{n+1}$. Each subsequent latent state $\boldsymbol{z}_{i}$ is decoded independently to predict the output $\hat{y}_{i}$ for $i>n$. This process is mathematically represented as follows:
@@ III-B Architecture-aware Differential Equation
We implement this using an RNN with GRU units for the sequence encoder. Other encoder options include Self-Attention (Vaswani et al., 2017) and Temporal Convolutional Networks (TCN) (Pandey and Wang, 2019), which adapt well to varying observation lengths. We conduct an ablation study in the experimental section to evaluate the performance of these alternative sequence encoder implementations.
@@
Assuming the graph representation $G$ contains $N$ nodes, the adjacency matrix $A\in\mathbb{R}_{\geq 0}^{N\times N}$ details node connections, where $A_{ij}>0$ represents the edge from node $i$ to node $j$. We analyze two foundational neural network types: MLPs and CNNs. In MLPs, nodes correspond to neurons and edges correspond to the presence of the connections between neurons, whereas in CNNs, nodes represent feature maps and edges depict operations like $1\times 1$ and $3\times 3$ convolutions, or $3\times 3$ average pooling. For CNNs, we adopt the cell-based representation (Dong and Yang, 2020; Liu et al., 2018), which consists of four principal building blocks: stems, normal cells, reduction cells, and classification heads. The stem block is a fixed sequence of convolutional layers to process the input images. This is followed typically by 14-20 cells with reduction cells placed at 1/3 and 2/3 of the total depth. A normal cell contains 4 nodes, each of which belongs to the set of operations: skip connections, identity or zero (indicating the presence or absence of connections between certain layers), $1\times 1$ and $3\times 3$ convolutions, and $3\times 3$ average pooling. Finally, the classification head employs a global pooling layer followed by a single fully connected layer and returns the network’s output. Since the cell is repeated throughout this macro-skeleton, a CNN can be uniquely represented by its cell.
To derive a graph-level embedding, we first implement node-level message passing, followed by global pooling to extract a global embedding. Our method differs from existing architecture-based performance prediction approaches (Liu et al., 2018; Wen et al., 2020; Knyazev et al., 2021) by treating the node as a feature map rather than focusing on operations. Nodal features are calculated from the in-degree and out-degree of each node, normalized by the total number of edges. For MLPs, edge features are binary, while for CNN cells, they are integers representing operation types: $\{\text{zeroize}:0,\text{$1\times 1$ conv}:1,\text{$3\times 3$ conv}:2,\text{$3\times 3$ avg pooling}:3\}$. The nodal feature matrix is denoted as $X\in\mathbb{R}^{N\times d}$. The graph encoding process returns a vector representation of the architecture.
@@
For node-level message passing, we employ GCN layers and normalize the adjacency matrix to stabilize training, following  (Kipf and Welling, 2016).
@@ A GCN layer then transforms the nodal features into a hidden representation:
The transformation from discrete to continuous domain is predicated on the assumption that the step size, or learning rate, approaches zero, thereby approximating the time derivative of NN parameters $\Theta$ using $\frac{d\Theta}{dt}=-\frac{\partial\mathcal{L}}{\partial\Theta}$ (Su et al., 2016), where $\mathcal{L}$, $\Theta$ denotes the objective function and parameters from the source task. Assuming the parameters implicitly depends on time $t$, the time derivative of the training loss can be written as:
@@
Here, $\dot{\boldsymbol{z}}$ denotes the derivative of the latent state vector $\boldsymbol{z}$ with respect to time, driven by the function $f_{\theta_{2}}$, which takes as input both the latent state $\boldsymbol{z}$ and a graph-level embedding $\boldsymbol{z}_{G}$. This combination allows the model to simultaneously consider the dynamic properties of the learning curve and the static characteristics of the architecture, enhancing the predictive capability of the system. Eq (9) can be regarded as a single-agent representation of the dynamics induced by NN training, which involves multiple trajectories of neurons, edge weights, and loss. This reduced representation of coupled dynamical system has been explored in (Gao et al., 2016; Laurence et al., 2019) to study the tipping point of the original network dynamics. The difference from the prior dynamical system reduction approach is that the graph reduction mechanism is learnable so that the model can be adapted to unseen trajectories derived from different optimization trials.
@@ IV Experiments
We consider test loss and test accuracy curves for both MLP-based and CNN-based architectures. Specifically for MLPs, we use car and segment tabular data binary classification tasks from OpenML (Vanschoren et al., 2014) as source tasks. Following LCBench (Zimmer et al., 2021), we randomly generate training configurations that include variables such as the number of layers, number of hidden units, and learning rates. For each dataset, we conduct 550 optimization trials across 200 epochs. For CNN-based models, we employ the NAS-Bench-201 dataset (Dong and Yang, 2020), which provides comprehensive learning curves for each architecture over a span of 200 epochs across two image datasets: CIFAR-10, CIFAR-100 (Krizhevsky et al., 2009). We randomly select 5,000 architectures from CIFAR-10 and CIFAR-100 to form our dataset. Furthermore, we reserve 20% of all trials as the test set for each MLP source task and 25% for each CNN source task.
@@ IV-A Extrapolating Real-world Learning Curves
The goal of this experiment is to evaluate the LC-GODE model against established learning curve prediction methods using real-world benchmarks. We train our model separately on the test loss curves of each source task. The condition length is set to 10 epochs for all methods in this experiment. The instantiation of LC-GODE that we report features: (i) an architecture encoder that utilizes 2 layers and employs a learnable pooling technique, (ii) an observed time-series encoder implemented using GRU, (iii) an ODE function with a 2-layer MLP and integrated using the Runge-Kutta 4 method (Butcher, 1996). Further details on the evaluation metrics and the training settings can be found in the supplemental materials.
@@
We evaluate our model against six methods, including three Bayesian approaches and three general time-series prediction approaches. (i) LC-BNN: This method utilizes a Bayesian neural network to model the posterior distribution of future learning curves. The probability function is constructed from a combination of basis functions, following the approach described by Domhan et al. (2015). (ii) LC-PFN: A transformer-based model is trained on synthetic curves that are generated from a pre-defined prior. This method serves as an efficient alternative to traditional Markov Chain Monte-Carlo (MCMC) techniques for sampling from posterior distributions. (iii) VRNN: A probabilistic model utilizing random forests and Bayesian recurrent neural networks. (iv) LSTM: This Recurrent Neural Network captures sequential dependencies. Before the observation cutoff at $n$, input states are derived from actual observations. Post $n$, the model uses its own predictions from previous timestamps as inputs. (v) NODE: Latent Ordinary Differential Equations, focusing primarily on modeling the latent loss representation without incorporating architectural information. (vi) NSDE: Latent Stochastic Differential Equations, consisting of a drift term and a diffusion term. The drift term is the same as NODE, and the sequence encoder and decoder is the same as in our model. Both NODE and NSDE are trained using the variational framework.